OpenStudy (anonymous):

Solve. 7x = 2x – 35 Solve. 1/5(25-5a)=4-a

OpenStudy (triciaal):

7x = 2x – 35 variable to one side constant to the other whatever you do to one side of an equation also do to the other to keep the balance 7x -2x = -35 and simplify if possible 5x = -35 divide by the coefficient of the variable (to get 1* variable) 5x/5 = -35/5 x = -7 check left side 7(-7)= -49 right side 2(-7)-35 = -7-35 = -49 left = right therefore yes x = -7

OpenStudy (triciaal):

now try the other one first multiply every term by the lowest common multiple to get rid of the fraction

OpenStudy (anonymous):

125-25a=4-a right?

OpenStudy (triciaal):

no every term is multiplied when you get rid of fractions actually you can factor out 5 on the left side

OpenStudy (anonymous):

I'm horrible with fractions :/

OpenStudy (triciaal):